# CHIPS: Efficient CLIP Adaptation via Curvature-aware Hybrid Influence-based Data Selection

### v1

@@ Abstract
Adapting CLIP to vertical domains is typically approached by novel fine-tuning strategies or by continual pre-training (CPT) on large domain-specific datasets. Yet, data itself remains an underexplored factor in this process. We revisit this task from a data-centric perspective: Can effective data selection substitute for large-scale datasets in CPT? We introduce CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), which assigns each image–text pair a utility score that integrates three complementary factors aligned with three goals: faithfulness via a curvature-aware, Newton-style alignment computed in CLIP’s end-point subspace; scalability via an InfoNCE-aware curvature estimator with Johnson–Lindenstrauss (JL) sketching; and retention via a selection-aware relevance weight combined with learnability to balance target adaptation against general-domain preservation. We justify this design theoretically by proving a lower‑bound guarantee on the proxy’s correlation with full‑parameter alignment and by characterizing the bias–variance trade‑offs introduced by curvature mixing and JL sketching. We evaluate CHIPS empirically across various settings: 1) CHIPS attains state-of-the-art performance among selection baselines on 17 medical benchmarks, matches full-dataset CPT with 30% of the data, and outperforms half-dataset CPT using only 10%; 2) on 31 general-domain benchmarks, CHIPS yields the smallest performance drop under 10–30% data-retention budgets. Code, data, and checkpoints will be released.
@@ 1 Introduction
Vision–language models such as CLIP [44] achieve strong zero-shot recognition in general domains, but their performance degrades sharply in vertical settings (e.g., medical imaging and biology) where vocabulary, acquisition protocols, and label taxonomies shift substantially [66]. To adapt CLIP to such domains, two main paradigms have emerged: 1) Model-centric methods modify training or parameterization (e.g., probabilistic fine-tuning [25], many-to-many contrastive learning [24], and PEFT variants [18, 46, 65, 58, 35]); 2) Data-centric methods pursue continual pre-training (CPT) on large, domain-specific datasets, ranging from millions to hundreds of millions of pairs across medical and biodiversity settings [34, 64, 36, 23, 48, 59, 50, 17, 60, 15, 54, 4]. However, for vertical domains, collecting, curating, and processing ever-larger datasets is costly, and indiscriminate upsampling can even dilute learning with redundant, low-utility samples. This naturally raises a question: Is sheer scale truly necessary for effective CPT?
@@
Meanwhile, a complementary line of work studies data attribution [9] to quantify how individual samples affect training dynamics and downstream generalization. Classical influence functions [28] and variants such as TracIn [43], EL2N [42], Arnoldi/GEX/FVM [47, 27, 62], were developed for single-tower model on supervised classification with additive cross-entropy and full-parameter updates. In CLIP, three mechanisms break these assumptions and systematically mis-rank examples. (A) Cross-modal curvature of dual encoders. CLIP’s dual encoders induce non-block-diagonal second-order curvature and block-diagonal proxies ignore this coupling and mis-rank samples. (B) Non-local gradients under InfoNCE. Each sample’s gradient depends on the softmax normalizer over the full negative set, making influence batch-/global-dependent rather than per-example additive and thus sensitive to batch size, queue design, and negative composition. (C) Endpoint dominance at the projection heads. Empirically, the projection heads and temperature drive early shifts in similarity distributions, whereas backbone updates are smaller and slower, rendering full-parameter updates partially unnecessary at the early stage.
@@
Motivated by (A)–(C), we propose CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), a CLIP-specific selector that scores each candidate by the expected one-step drop of the target evaluation loss and selects the top examples for CPT. CHIPS is designed for three goals, faithfulness, scalability, and retention, with one component per goal: (i) a Newton-style alignment computed on the projection heads and temperature that provably lower-bounds full-parameter alignment and enjoys better conditioning (Sec. 2.2); (ii) an InfoNCE-aware curvature preconditioner built from positive/negative gradient moments and compressed via Johnson–Lindenstrauss (JL) sketch to achieve near-linear time/memory with a quantified variance–bias trade-off (Sec. 2.3); and (iii) a selection-aware soft weighting that combines per-sample learnability with target-domain relevance to explicitly control the adaptation–retention balance (Sec. 2.4). Tab. 1 summarizes the dimensions that matter in practice and positions CHIPS against several common baselines.
@@
Across 17 medical benchmarks, CHIPS matches full-dataset CPT using only 30% of the training pool, and outperforms half-dataset CPT using merely 10% data, achieving state-of-the-art among selection baselines. For 31 general-domain benchmarks, although domain adaptation inevitably reduces average performance, CHIPS consistently reduces the drop relative to the previous SOTA selector, indicating stronger retention of general-domain capabilities. In short, with principled attribution and CLIP-aware curvature, effective CPT does not require extreme scale.
Our contributions are threefold:
  • Curvature-aware proxy alignment in CLIP’s end-point subspace. We define a Newton-inspired alignment that ranks samples by expected one-step decrease on the target validation loss and provably lower-bounds full-parameter alignment.
  • InfoNCE-aware curvature estimation with JL sketching and guarantees. We incorporate negative-pair geometry to capture cross-example curvature and characterize the variance–bias trade-off of the estimator.
  • Selection-aware weighting for learnability and retention. We combine learnability and domain-relevance weights to emphasize decision-boundary, on-domain samples while softly limiting selection drift, mitigating catastrophic forgetting of general-domain knowledge.
@@ 2 Method
CHIPS ranks training samples for CPT by estimating, for each sample, how much a one-step update would decrease the evaluation loss on the target domain. We operationalize this idea along three tightly coupled components: (i) a curvature-aware proxy alignment score computed in CLIP’s end-point geometry (projection heads and temperature), (ii) a negative-pair curvature estimator that captures the second-order coupling induced by symmetric InfoNCE and scales via JL sketching, and (iii) two multiplicative weights for learnability and target-domain relevance, forming a single selection utility. Fig. 1 provides the workflow of CHIPS.
@@ 2.1 Problem Setup
We consider a pre-trained general CLIP model with parameters $\theta$ and a large target-domain training pool $\mathcal{D}_{\text{pool}}=\{z_{i}\}_{i=1}^{N}$ of image–text pairs. Our goal is to select a valuable subset from $\mathcal{D}_{\text{pool}}$ for CPT so as to maximize performance on a held-out target-domain evaluation set $\mathcal{D}_{\text{eval}}$ consisting of samples from various validation sets of downstream tasks. Let $\mathcal{L}_{\text{eval}}(\theta)\triangleq\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\ell(z;\theta)]$ denote the evaluation loss, where $\ell$ is the symmetric InfoNCE objective [44] with a learnable logit-scale (temperature) $\tau>0$. Denote the evaluation mean gradient by $\mathbf{u}\triangleq\nabla_{\theta}\mathcal{L}_{\text{eval}}(\theta)=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\nabla_{\theta}\ell(z;\theta)]$. During CPT, one Stochastic Gradient Descent (SGD) step draws a mini-batch of size $B$ from a selection distribution $q$ and computes the batch gradient $\widehat{\mathbf{g}}=\tfrac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\ell(z_{i};\theta)$ to update parameters via
$\theta^{\prime}=\theta-\eta\,\widehat{\mathbf{g}}$, (1)
where $\eta>0$ is the learning rate. We also define the training mean gradient under $q$ by $\mathbf{g}_{q}\triangleq\mathbb{E}_{z\sim q}[\nabla_{\theta}\ell(z;\theta)]$. Assume $\mathcal{L}_{\text{eval}}$ is twice continuously differentiable in a neighborhood of $\theta$ and let $\mathbf{H}_{\text{eval}}(\theta)=\nabla_{\theta}^{2}\mathcal{L}_{\text{eval}}(\theta)$. A second-order Taylor expansion with integral remainder and Hessian-Lipschitz constant $L_{H}$ yields, for $\Delta\theta=-\eta\,\widehat{\mathbf{g}}$,
@@ 2.2 Curvature-aware Proxy Alignment
Given CLIP’s dual-encoder with projection heads and a temperature, we implement CHIPS in the end-point subspace consisting of projection heads and temperature $\vartheta=\{\mathbf{W}_{v},\mathbf{W}_{t},\tau\}$. For a sample $z$, let $\mathbf{g}_{\vartheta}(z)=\nabla_{\vartheta}\ell(z;\theta)$ be its gradient restricted to $\vartheta$, and define the evaluation mean $\mathbf{u}_{\vartheta}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\vartheta}(z)]$. For reference, denote the full-parameter gradient by $\mathbf{g}_{\theta}(z)=\nabla_{\theta}\ell(z;\theta)$ and its evaluation mean by $\mathbf{u}=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\mathbf{g}_{\theta}(z)]$. Motivated by the one-step descent view in Sec. 2.1, an ideal update direction on $\vartheta$ is the Newton direction $\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$, where $\mathbf{H}_{\vartheta}$ denotes a population curvature (e.g., generalized Gauss–Newton) on the subspace. We therefore define a curvature-aware proxy Newton alignment for each sample:
@@
where $\mathbf{M}\succ 0$ is a tractable curvature surrogate instantiated in Sec. 2.3. The sign and magnitude of $A(z)$ indicate whether (and how strongly) a one-step update on $z$ moves the model along a descent direction for the evaluation loss: large positive $A(z)$ is preferred.
Directly computing alignment with respect to all parameters is expensive and brittle to high-dimensional noise. We instead argue that the end-point geometry is a reliable proxy for full alignment. Eq. (4) can be rewritten as $A(z)=\tilde{\mathbf{g}}_{\vartheta}(z)^{\top}\tilde{\mathbf{u}}_{\vartheta}$ with $\tilde{\mathbf{g}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{g}_{\vartheta}$ and $\tilde{\mathbf{u}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{u}_{\vartheta}$. Thus $\mathbf{M}$ is a fixed (sample-independent) preconditioner and the proxy–full correlation bound applies unchanged after the congruent transform, i.e., replace $\Sigma_{g}$ by $\mathbf{M}^{-1/2}\Sigma_{g}\mathbf{M}^{-1/2}$ and $\mathbf{S}$ by $\mathbf{M}^{1/2}\mathbf{S}\,\mathbf{M}^{1/2}$. We can state a single unified bound below.
Consider the following local linearization:
@@
and define $X(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{u}_{\vartheta}$, $Y(z)=\mathbf{g}_{\theta}(z)^{\top}\mathbf{u}$ with Pearson correlation $\rho_{XY}$, $\Sigma_{g}=\operatorname{Cov}[\mathbf{g}_{\vartheta}(z)]$, $\mathbf{S}=\tfrac{1}{2}(\mathbf{J}^{\top}\bar{\mathbf{J}}+\bar{\mathbf{J}}^{\top}\mathbf{J})$, $\mathbf{B}=\Sigma_{g}^{1/2}\mathbf{S}\,\Sigma_{g}^{-1/2}$, and $\sigma_{\zeta}^{2}=\operatorname{Var}[\zeta(z)]$ for the mismatch term induced by $(\mathbf{r},\boldsymbol{\varepsilon})$.
Theorem 1 (Proxy–full alignment correlation).
If $\zeta(z)$ is uncorrelated with $\mathbf{g}_{\vartheta}(z)$, then
@@
The matrices $\mathbf{S}$ and $\Sigma_{g}$ summarize how end-point gradients relate to full-parameter gradients and how they vary. When $\mathbf{S}$ is well-conditioned along $\mathrm{span}(\mathbf{u}_{\vartheta})$ and the mismatch variance $\sigma_{\zeta}^{2}$ is modest, the bound keeps $\rho_{XY}$ bounded away from zero, so ranking by the proxy alignment tracks the full alignment. Under approximate isotropy or commuting structure, the expression simplifies by effectively replacing $\mathbf{B}$ with $\mathbf{S}$ (see App. B.2). Empirically, on a 100K subset of BIOMEDICA [36], Spearman’s $\rho_{s}=0.83$ between proxy and full alignment supports this proxy view.
@@ 2.3 Curvature Estimation
Symmetric InfoNCE couples each positive pair with many negatives through the softmax normalizer, creating cross-example curvature that a positive-only or diagonal surrogate fails to capture. Since our alignment score $A(z)$ relies on a curvature-preconditioned inner product, missing this cross mass yields a biased Newton proxy and unstable rankings along highly coupled directions. Working in the end-point subspace $\vartheta$, we approximate the population curvature (e.g., generalized Gauss–Newton) by combining self and cross second moments of subspace gradients:
@@
with a mixing weight $\alpha\in[0,1]$ and a small Tikhonov parameter $\lambda>0$ to ensure numerical stability in matrix inversion. Plugging $\mathbf{M}$ into Eq. (4) yields the curvature-aware proxy Newton score $A(z)$. By mixing $\Phi_{\mathrm{pos}}$ with the cross moment $\Phi_{\mathrm{neg}}$ (with $\alpha>0$), we restore the off-diagonal curvature induced by InfoNCE’s negatives that positive-only surrogates ignore. The surrogate $\mathbf{M}\!\succ\!0$ thus preconditions $A(z)$ in a fixed metric without altering the proxy–full correlation framework (Sec. 2.2); $\alpha$ controls coupling strength and $\lambda$ provides ridge stability. This yields an alignment score that is sensitive to the true second-order structure yet robust. Let the ideal scalar be $A^{\star}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$ and the curvature-approximated score $A_{\alpha}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$ with $\mathbf{M}$ as in Eq. (8), and define $\Delta_{\alpha}=\mathbf{H}_{\vartheta}-\mathbf{H}_{\vartheta}^{(\alpha)}$. For further efficiency, we apply a $k$-dimensional JL sketch to compute a sketched score $\widehat{A}_{\alpha}(z)$. We discuss the choice of JL sketch in Sec. 4.1.

### v2

@@ Abstract
Adapting CLIP to vertical domains is typically approached by novel fine-tuning strategies or by continual pre-training (CPT) on large domain-specific datasets. Yet, data itself remains an underexplored factor in this process. We revisit this task from a data-centric perspective: Can effective data selection substitute for large-scale datasets in CPT? We introduce CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), which assigns each image-text pair a utility score that integrates three complementary factors aligned with three goals: faithfulness via a curvature-aware and Newton-style alignment computed in CLIP’s end-point subspace; scalability via an InfoNCE-aware curvature estimator with Johnson-Lindenstrauss (JL) sketching; and retention via a selection-aware relevance weight combined with learnability to balance target adaptation against general-domain preservation. We justify this design theoretically by proving a lower-bound guarantee on the proxy’s correlation with full-parameter alignment and by characterizing the bias-variance trade-offs introduced by curvature mixing and JL sketching. We evaluate CHIPS empirically across various settings: 1) CHIPS attains state-of-the-art performance among selection baselines on 17 medical benchmarks, matches full-dataset CPT with 30% of the data, and outperforms half-dataset CPT using only 10%; 2) on 31 general-domain benchmarks, CHIPS yields the least performance drop under all retention ratios.
@@ 1 Introduction
Vision-language models such as CLIP [44] achieve strong zero-shot recognition in general domains, but their performance degrades sharply in vertical settings (e.g., medical imaging and biology) where vocabulary, acquisition protocols, and label taxonomies shift substantially [66]. To adapt CLIP to such domains, two main paradigms have emerged: 1) Model-centric methods modify training or parameterization (e.g., probabilistic fine-tuning [25], many-to-many contrastive learning [24], and PEFT variants [18, 46, 65, 58, 35]); 2) Data-centric methods pursue continual pre-training (CPT) on large, domain-specific datasets, ranging from millions to hundreds of millions of pairs across medical and biodiversity settings [34, 64, 36, 23, 48, 59, 50, 17, 60, 15, 54, 4]. However, for vertical domains, collecting, curating, and processing increasingly large datasets is costly, and indiscriminate upsampling can even impair learning by introducing redundant, low-utility samples. This naturally raises a question: Is sheer scale truly necessary for effective CPT?
@@
Meanwhile, a complementary line of work studies data attribution [9] to quantify how individual samples affect training dynamics and downstream generalization. Classical influence functions [28] and variants such as TracIn [43], EL2N [42], Arnoldi/GEX/FVM [47, 27, 62], were developed for single-tower model on supervised classification with additive cross-entropy and full-parameter updates. In CLIP, three mechanisms break these assumptions and systematically mis-rank examples. (A) Cross-modal curvature of dual encoders. CLIP’s dual encoders induce non-block-diagonal second-order curvature and block-diagonal proxies ignore this coupling and mis-rank samples. (B) Non-local gradients under InfoNCE. Each sample’s gradient depends on the softmax normalizer over the full negative set, making influence batch-/global-dependent rather than per-example additive and thus sensitive to batch size, queue design, and negative composition. (C) Endpoint dominance at the projection heads. Empirically, the projection heads and temperature parameter drive early shifts in similarity distributions, rendering full-parameter influence computation relatively unnecessary for CLIP.
@@
Motivated by (A)-(C), we propose CHIPS (Curvature-aware Hybrid Influence in Projection Subspace), a CLIP-specific selector that scores each candidate by the expected one-step drop of the target evaluation loss and selects the top examples for CPT. CHIPS is designed for three goals, faithfulness, scalability, and retention, with one component per goal: (i) a Newton-style alignment computed on the projection heads and temperature that provably lower-bounds full-parameter alignment and enjoys better conditioning (Sec. 2.2); (ii) an InfoNCE-aware curvature preconditioner built from positive/negative gradient moments and compressed via Johnson-Lindenstrauss (JL) sketching to achieve near-linear time/memory with a quantified variance-bias trade-off (Sec. 2.3); and (iii) a selection-aware soft weighting that combines per-sample learnability with target-domain relevance to explicitly control the adaptation-retention balance (Sec. 2.4). Tab. 1 positions CHIPS against several common baselines from practical perspectives.
@@
Across 17 medical benchmarks, CHIPS matches full-dataset CPT using only 30% of the training pool, and outperforms half-dataset CPT using merely 10% data, achieving SOTA among selection baselines. For 31 general-domain benchmarks, while domain adaptation inevitably reduces average performance, CHIPS consistently reduces the drop relative to the previous SOTA selector, indicating stronger retention of general-domain capabilities. In short, with principled attribution and CLIP-aware curvature, effective CPT does not require extreme scale.
Our contributions are threefold [^1] https://github.com/mihara-bot/CHIPS.. First, we propose an alignment score computed in end-point space of CHIP, which ranks samples by expected one-step decrease on evaluation loss. Second, we incorporate negative-pair geometry to capture cross-example curvature and analyze the variance-bias trade-off of the estimator. Third, we combine learnability and domain-relevance weights to emphasize decision-boundary, on-domain samples while softly limiting selection drift and mitigating catastrophic forgetting. Experimental results demonstrate the effectiveness of CHIPS on both medical and general downstream tasks.
@@ 2 Method
CHIPS ranks training samples using an alignment score that estimates how much a one-step update on a given sample would reduce the evaluation loss on the target domain, together with auxiliary learnability and domain relevance scores. More specifically, CHIPS consists of three tightly coupled components: (i) a curvature-aware proxy alignment score computed in CLIP’s end-point geometry, namely the projection heads and temperature; (ii) a negative-pair curvature estimator that captures the second-order coupling induced by symmetric InfoNCE and is made scalable through JL sketching; and (iii) two multiplicative weights for learnability and target-domain relevance, forming a single selection utility. Fig. 1 provides the workflow of CHIPS.
@@ 2.1 Problem Setup
We consider a pre-trained general CLIP model with parameters $\theta$ and a large target-domain training pool $\mathcal{D}_{\text{pool}}=\{z_{i}\}_{i=1}^{N}$ of image-text pairs. Our goal is to select a valuable subset from $\mathcal{D}_{\text{pool}}$ for CPT so as to maximize performance on a held-out target-domain evaluation set $\mathcal{D}_{\text{eval}}$ consisting of samples from various validation sets of downstream tasks. Let $\mathcal{L}_{\text{eval}}(\theta)\triangleq\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\ell(z;\theta)]$ denote the evaluation loss, where $\ell$ is the symmetric InfoNCE objective [44] with a learnable logit-scale (temperature) $\tau>0$. Denote the evaluation mean gradient by $\mathbf{u}\triangleq\nabla_{\theta}\mathcal{L}_{\text{eval}}(\theta)=\mathbb{E}_{z\sim\mathcal{D}_{\text{eval}}}[\nabla_{\theta}\ell(z;\theta)]$. During CPT, one Stochastic Gradient Descent (SGD) step draws a mini-batch of size $B$ from a selection distribution $q$ and computes the training gradient $\widehat{\mathbf{g}}=\tfrac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\ell(z_{i};\theta)$ to update parameters via $\theta^{\prime}=\theta-\eta\,\widehat{\mathbf{g}}$, where $\eta>0$ is the learning rate. We also define the training mean gradient under $q$ by $\mathbf{g}_{q}\triangleq\mathbb{E}_{z\sim q}[\nabla_{\theta}\ell(z;\theta)]$. Assume $\mathcal{L}_{\text{eval}}$ is twice continuously differentiable in a neighborhood of $\theta$ and let $\mathbf{H}_{\text{eval}}(\theta)=\nabla_{\theta}^{2}\mathcal{L}_{\text{eval}}(\theta)$. A second-order Taylor expansion with integral remainder and Hessian-Lipschitz constant $L_{H}$ yields the following bound for $\Delta\theta=-\eta\,\widehat{\mathbf{g}}$:
@@ 2.2 Curvature-aware Proxy Alignment
CHIPS is implemented on the end-point subspace of CLIP $\vartheta=\{\mathbf{W}_{v},\mathbf{W}_{t},\tau\}$, consisting of the two projection heads and the temperature. For a sample $z$, let $\mathbf{g}_{\vartheta}(z)=\nabla_{\vartheta}\ell(z;\theta)$ denote its gradient restricted to $\vartheta$, and let $\mathbf{u}_{\vartheta}$ be the corresponding mean evaluation gradient. Motivated by the one-step descent view in Sec. 2.1, the ideal update direction on this subspace is the Newton direction $\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$, where $\mathbf{H}_{\vartheta}$ denotes a population curvature on $\vartheta$. Accordingly, we define the curvature-aware proxy alignment score:
@@
where $\mathbf{M}\succ 0$ is a tractable curvature surrogate detailed in Sec. 2.3. A large positive $A(z)$ means that a one-step update on $z$ moves the model more strongly along a descent direction for the evaluation loss.
Our central hypothesis is that alignment measured in the end-point subspace is a reliable proxy for alignment in the full parameter space. Eq. (3) can be rewritten as $A(z)=\tilde{\mathbf{g}}_{\vartheta}(z)^{\top}\tilde{\mathbf{u}}_{\vartheta}$ with $\tilde{\mathbf{g}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{g}_{\vartheta}$ and $\tilde{\mathbf{u}}_{\vartheta}=\mathbf{M}^{-1/2}\mathbf{u}_{\vartheta}$. This indicates that the curvature-aware score is simply the standard inner-product alignment after a fixed, sample-independent preconditioning. Therefore, the proxy-full analysis below applies directly to this curvature-aware form. The explicit transformed covariance and coupling operators are given in App. B.2. To formalize this proxy view, consider the local linearization
@@
where $\mathbf{J}$ is a local linear map that lifts a sample’s end-point gradient $\mathbf{g}_{\vartheta}(z)$ to its full-parameter gradient, and $\bar{\mathbf{J}}$ is the analogous local linear map for the mean evaluation gradient. The residual terms $\mathbf{r}(z)$ and $\boldsymbol{\varepsilon}$ capture the corresponding approximation errors.
Define the subspace alignment $X(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{u}_{\vartheta}$ and the full alignment $Y(z)=\nabla_{\theta}\ell(z)^{\top}\mathbf{u}$. Let $\Sigma_{g}=\operatorname{Cov}[\mathbf{g}_{\vartheta}(z)]$, $\mathbf{S}=\tfrac{1}{2}(\mathbf{J}^{\top}\bar{\mathbf{J}}+\bar{\mathbf{J}}^{\top}\mathbf{J})$, $\mathbf{B}=\Sigma_{g}^{1/2}\mathbf{S}\,\Sigma_{g}^{-1/2}$, and let $\sigma_{\zeta}^{2}=\operatorname{Var}[\zeta(z)]$ denote the variance of the mismatch term induced by $(\mathbf{r},\boldsymbol{\varepsilon})$. We have the following lower bound:
Theorem 1 (Proxy-full alignment correlation).
If $\zeta(z)$ is uncorrelated with $\mathbf{g}_{\vartheta}(z)$, then the Pearson correlation between $X(z)$ and $Y(z)$ satisfies
@@
Theorem 1 shows that the proxy score tracks the full score when two conditions hold: (i) the subspace-to-full map is well aligned along $\mathbf{u}_{\vartheta}$, as summarized by $\mathbf{S}$, and (ii) the residual mismatch variance $\sigma_{\zeta}^{2}$ is not too large. Under approximate isotropy or commuting structure, the bound simplifies further (provided in App. B.2).
However, CHIPS uses the alignment score for ranking samples. Therefore, the relevant empirical quantity is the Spearman correlation between proxy and full scores. On a 100K subset of BIOMEDICA [36], we observe a Spearman correlation of 0.83, proving that end-point subspace scores preserve the ordering induced by full alignment well.
@@ 2.3 Curvature Estimation
Symmetric InfoNCE couples each positive pair with multiple negatives through the softmax normalizer, creating cross-example curvature that a positive-only or diagonal surrogate fails to capture. Since our alignment score $A(z)$ relies on a curvature-preconditioned inner product, missing this cross mass yields a biased Newton proxy and unstable rankings along highly coupled directions. Working in the end-point subspace $\vartheta$, we approximate the population curvature by combining self and cross second moments of subspace gradients:
@@
with a mixing weight $\alpha\in[0,1]$ and a small Tikhonov parameter $\lambda>0$ to ensure numerical stability in matrix inversion. Plugging $\mathbf{M}$ into Eq. (3) yields the curvature-aware proxy Newton score $A(z)$. By mixing $\Phi_{\mathrm{pos}}$ with the cross moment $\Phi_{\mathrm{neg}}$ (with $\alpha>0$), we restore the off-diagonal curvature induced by InfoNCE’s negatives that positive-only surrogates ignore. The surrogate $\mathbf{M}\!\succ\!0$ thus preconditions $A(z)$ in a fixed metric without altering the proxy-full correlation framework (Sec. 2.2); $\alpha$ controls coupling strength and $\lambda$ provides ridge stability. This yields an alignment score that is sensitive to the true second-order structure yet robust. Let the ideal scalar be $A^{\star}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{H}_{\vartheta}^{-1}\mathbf{u}_{\vartheta}$ and the curvature-approximated score $A_{\alpha}(z)=\mathbf{g}_{\vartheta}(z)^{\top}\mathbf{M}^{-1}\mathbf{u}_{\vartheta}$ with $\mathbf{M}$ as in Eq. (7), and define $\Delta_{\alpha}=\mathbf{H}_{\vartheta}-\mathbf{H}_{\vartheta}^{(\alpha)}$. For further efficiency, we apply a $k$-dimensional JL sketch to compute a sketched score $\widehat{A}_{\alpha}(z)$. We discuss the choice of JL sketch in Sec. 4.1.